### v2

# ]The authors are members of LabEx Tec21 (Investissements d’Avenir - grant agreement ANR-11-LABX-0030)

Abstract
Blood rheology and microcirculation are strongly influenced by red blood cell (RBC) aggregation. The aggregability of RBCs can vary significantly due to factors such as their mechanical and membrane surface properties, which are affected by cell aging in vivo. In this study, we investigate RBC aggregability as a function of their density, a marker of cell age and mechanical properties, by separating RBCs from healthy donors into different density fractions using Percoll density gradient centrifugation. We examine the dissociation rates of aggregates in a controlled medium supplemented with Dextran, employing an extensional flow technique based on hyperbolic microfluidic constrictions and image analysis, assisted by a convolutional neural network (CNN). In contrast to other techniques, our microfluidic experimental approach highlights the behavior of RBC aggregates in dynamic flow conditions relevant to microcirculation. Our results demonstrate that aggregate dissociation is strongly correlated with cell density and that aggregates formed from the denser fractions of RBCs are significantly more robust than those from the average cell population. This study provides insight into the effect of RBC aging in vivo on their mechanical properties and aggregability, underscoring the importance of further exploration of RBC aggregation in the context of cellular senescence and its potential implications for hemodynamics. Additionally, it suggests that this technique can complement existing methods for improved evaluation of RBC aggregability in health and disease.
Blood, red blood cell, aggregation, RBC aggregability, RBC aging, RBC density, Percoll, artificial intelligence, CNN
preprint: APS/123-QED
I Introduction
Red blood cell (RBC) aggregation is a critical factor influencing blood rheology and circulation. Due to the presence of plasma proteins, RBCs tend to reversibly aggregate, forming mostly linear structures known as rouleaux favored by their flat biconcave shape. These linear structures can assemble into larger, gel-like networks at low shear rates and are mostly responsible for the strong shear-thinning of blood observed at low shear rates chien1970 , in conjunction with RBC deformability chien1967a ; lanotte16 .
Although shear rates in blood circulation generally exceed 100 s-1 Robertson2008 , where aggregation does not impact macroscopic rheology measurements chien1970 , various studies have shown that aggregation interactions play a crucial role in shaping the structure of RBC suspensions in narrow channels Bishop2001 ; Zhang2009 , stabilizing clusters within capillaries brust14 , and affecting blood perfusion in vascular networks Reinhart2017 . Indeed, in narrow vessels or channels, small RBC aggregates made up of a few cells can resist to strain rates up to a few 100 s-1 puthumana2024dissociation .
RBC aggregation is caused by plasma proteins, especially fibrinogen baskurt2011 , through two proposed mechanisms, both of which being supported by evidence. The first one is a bridging mechanism in which macromolecules such as fibrinogen adsorb (specifically or not) onto RBC membranes and form bridges between adjacent cells chien1987physicochemical ; Chien1973RedCA . The second mechanism is of purely entropic nature and involves depletion forces asakura1958 related to the finite size of macromolecules and was quantitatively predicted by Neu and Meiselmann neu2002 . While variations in plasma composition have an impact on aggregation, e.g. through fibrinogen production in inflammatory conditions castell1990acute which is usually reflected in erythrocyte sedimentation rate (ESR) tests bedell1985 , cellular factors also strongly influence the ability of cells to aggregate (usually referred to as aggregability baskurt2013erythrocyte ) and the properties of these aggregates. For instance, previous modelling and experimental studies have shown that cellular parameters such as cell volume or membrane bending and shear moduli have a significant impact on the equilibrium shape of erythrocyte aggregates flormann2017 ; hoore2018 . RBCs under normal conditions are highly deformable. Under pathological conditions, such as sickle cell anemia and bacterial infections, deformability of RBCs is greatly reduced. Clinical studies have shown that under such conditions aggregability of RBCs is also increased tripette2009red ; Baskurt1997Red ; Chen1996Enhanced ; ChongMartinez2003 . However reduced deformability is not the only factor leading to increased aggregability of RBCs and modifications of membrane surface properties such as charge or protein distribution also lead to major changes in aggregability via changes in electrostatic or binding interactions. For instance it was shown that elevated concentrations of plasma sodium lead to impairment of the erythrocyte glycocalyx and alteration of the zeta-potential, resulting in higher aggregability McNally2023 ; Oberleithner2013 while on the other hand experiments with artificially rigidified cells suggest that a small fraction of these cells mixed with a healthy sample tends to reduce the aggegation index and mean aggregate size kuck2022impact .
Even in healthy conditions, RBC properties are significantly heterogeneous and evolve in time. Produced in bone marrow, RBCs have a lifespan of about 120 days during which aging leads to increased density, stiffness and other alterations Waugh1992Mar ; bosch1994 ; franco2012measurement ; franco2013 before they are eliminated by phagocytosis by macrophages in the spleen. As blood flows through the circulatory system, the stresses experienced by blood components vary as they navigate the complex geometries and varying diameters of the circulatory system caro2012mechanics ; nichols2005mcdonald . These flow events can cause RBCs to elongate, alter membrane protein distribution, and potentially lead to membrane fatigue, impacting cell functionality and lifespan lipowsky2005microvascular ; pivkin2008accurate . As the cells age, they undergo several physical and chemical changes including loss of constituents like - water,(2,3)- bisphosphosphoglyceric acid, ATP, proteins, Hb, sialoglycoproteinssialic acids (SA), resulting structural changes to the cell membrane huang2011human ; lutz1983senescent ; lutz1987naturally .
There is no consensus on a strict equivalence between red cell density and age and studies show that there is a wide distribution of densities among cells of the same age, it is still accepted that on average density increases with cell age morrison1983does . However, a robust feature is that cell density is related to hemoglobin content mohandas86scattering , which has a direct and strong impact on cytoplasm viscosity. In addition, density increase was found to be correlated to loss of membrane surface area. Both phenomena lead to a decrease of static deformability (deformability index) and dynamic deformability (increase of relaxation time). Besides, density increase is also observed in RBC disorders associated with decreased cell deformability mohandas88densdistrib .
Recently, several studies have used Percoll or iodixanol density gradients to separate RBCs by density and analyze their dynamics, demonstrating a significant impact on RBC shape in flow nouaman2023effect , as well as migration and axial dispersion in capillaries losserand2019migration ; losserand2023axial .
Various techniques have been proposed to quantify aggregation and characterize the aggregability of erythrocytes at the suspension scale. These include the classic erythrocyte sedimentation rate (ESR) test bedell1985 , ektacytometry dacosta2016 , LORCA hardeman1994laser , and ultrasound scattering franceschini2010ultrasound , all of which are applied to dense suspensions. At the cellular scale, force measurements or controlled force application methods, such as micropipette aspiration techniques tozeren1989theoretical ; buxbaum1982quantitation , AFM steffen2013quantification , and more recently optical tweezers (OT) lee2017assessment ; Ermolinskiy2020Oct , allow direct investigation of cell-cell interactions. Notably, Ermolinskiy et al. Ermolinskiy2020Oct showed that the aggregation force measured by OT increases with the density of erythrocytes, both in autologous plasma and in 50 mg/ml Dextran 70 kDa solutions, suggesting an increase of RBC aggregability with cell age.
Recently, a new microfluidic technique combined with machine learning analysis was proposed to investigate the dissociation of small RBC aggregates in confined elongational flows relevant to the hydrodynamic stresses experienced in microcirculatory networks puthumana2024dissociation . This high-throughput technique provides detailed information about dissociation statistics as a function of strain rate for aggregates of 2 to 4 cells over large samples, complementing other techniques.
In this work, we use this technique to quantify the variations in dissociation probabilities of small aggregates of healthy erythrocytes in extensional flow as a function of cell density after separating them using Percoll density gradients. We show that aggregates formed from the denser fractions of red blood cells are significantly more robust than the average cell population and require significantly higher flow stresses for dissociation. These results suggest that in addition to decreased deformability, RBC aging leads to increased aggregation in microcirculatory conditions.
II Methods
II.1 Blood sample collection and preparation
Figure 1: RBC density fractionation using percoll solution gradient. a) 2 mL of 60% Hematocrit along with PBS is carefully placed above the $5\times 2$ mL Percoll density gradient solution. The densities of each layer is shown in the figure. b) Separated RBC layers, after centrifugation for 30 minutes at 4000 $\times$ g.
Blood samples were collected from three healthy male volunteers with informed consent through venipuncture, using tubes containing EDTA (1.6 mg/mL, SARSTEDT, Nümbrecht, Germany) as anticoagulant. Plasma and buffy coat were then separated from red blood cells by centrifugation (3000g, 9 minutes) prior to washing 3 times (3000g, 3 minutes) using PBS (Phosphate Buffered Saline, Gibco, pH 7.4, Fisher Scientific, Germany).
The separation of RBCs according to cell age was conducted by utilizing the density variations that arise during the ageing process Waugh1992Mar . To obtain density fractions of RBCs from a whole blood sample, we employed the Percoll density gradient centrifugation method outlined by Ermolinskiy et al. Ermolinskiy2020Oct . The Percoll solution (Cytiva 17-0891-01, Sigma-Aldrich, Taufkirchen, Germany) consists of Percoll, distilled water, and a 1.5M NaCl solution in different proportions, which creates the five different density levels for the experiment. Five solutions, with densities of 1.085 g/mL, 1.092 g/mL, 1.101 g/mL, 1.107 g/mL, and 1.122 g/mL, were prepared and carefully placed one over the other in a 15 mL centrifugation tube (highest density solution at the bottom, lowest at the top, with each layer of volume of 2 mL). Finally, a 2 mL suspension of washed red blood cells, diluted to a hematocrit of 60% in PBS, was carefully layered onto the Percoll gradient solution. The sample was then centrifuged at 4000 g for 30 minutes at a controlled temperature of 4°C. Centrifugation results in cells being fractionated and forming bands labelled $L_{0}-L_{4}$ in the transition zones between Percoll layers as shown in Fig. 1, with $L_{2}$ being the most concentrated layer i.e. corresponding to the peak of RBC density distribution in blood samples mohandas88densdistrib ; Ermolinskiy2020Oct ; nouaman2023effect . These different layers were then pipetted, washed, centrifuged and resuspended in a solution of 3 g/dL Dextran 70kDa to promote RBC aggregation with an interaction energy that is comparable to fibrinogen concentrations in the 4-5 g/L range neu2002 ; steffen2013quantification ; brust14 before using them separately in experiments. Due to the very low cell concentrations in layers $L_{0}$ and $L_{1}$, only layers $L_{2}$, $L_{3}$ and $L_{4}$ were used in experiments.
II.2 Microfluidics
Figure 2: Hyperbolic flow geometry producing extensional flow: (a) Design of the central part of the microfluidic channel and notations. Blue areas are image processing ROIs; (b) Example temporal sequence of a 2-cell aggregate breaking into $1+1$ cells under extensional flow. Time-step: 1.1 ms; Extension rate $\dot{\varepsilon}=$ 118 s-1.
The core component of the microfluidics setup, whose principle was described in detail in a previous study and was demonstrated to be a reliable tool to characterize RBC disaggregation puthumana2024dissociation , is a hyperbolic constriction geometry (Fig. 2 a). This design has been effectively implemented in prior studies bento2018deformation ; faustino2019microfluidic to induce an extensional flow thanks to a channel width that decreases inversely with the coordinate $x$ in the flow direction ($w(x)\sim 1/x$) and a constant channel thickness $h$. Mass conservation implies that the cross-sectional average of the axial velocity $<u_{x}(x)>$ increases linearly along $x$ i.e. the average extension rate $d<u_{x}(x)>/dx$ is constant. Particles flowing along the central line therefore experience an extensional strain rate $\dot{\varepsilon_{f}}={dv_{x}}/{dx}$ where $v_{x}=u_{x}(x,0,0)$. A series of such constrictions with constant thickness $h=20$ $\mu$m, in the $z$ direction, inlet width($w_{0}=200$ $\mu$m) and outlet width ($w_{L}=20$ $\mu$m) was implemented, with decreasing lengths, $L$ (1000 $\mu$m, 500 $\mu$m, 250 $\mu$m, 125  $\mu$m), inducing decreasing extension rates $\dot{\varepsilon_{f}}$ for a given flow rate. The measurements of aggregate dissociation statistics presented here were conducted in the hyperbolic channel with $L=250$ $\mu$ m as in puthumana2024dissociation .
The entrance section of the microfluidic chip is a 3-inlet flow-focusing device where the concentrated RBC suspension is focused in the $y$-direction between two perpendicular inlets of the suspending medium. It is followed by a $\sim$ 5 cm long channel, which allows RBC aggregates to relax and focus around the center plane in the $z$-direction thanks to migration forces losserand2019migration ; grandchamp13 ; podgorski11 . The flow is driven by applying pressure at 3 inlets of the microfluidics chip, using an Elveflow OB1 pressure controller.
[…]
II.3 Microscopy and Image processing
RBC cluster images were acquired using an inverted microscope (Eclipse TE2000-S, Nikon) equipped with a x20 objective lens and a high-speed camera (Fastec HiSpec 2G, FASTEC Imaging), providing a 2.05 pixels/$\mu$m resolution and a frame rates ranging 250-3000 fps. A typical aggregate dissociation sequence is shown in Fig. 2(b). The actual extension rate experienced by aggregates $\dot{\varepsilon}=dv_{p}/dx$ where $v_{p}$ is the particle velocity was evaluated for each value of the inlet pressures by measuring sequential positions $x(t)$ of a few particles (cells or aggregates) along the hyperbolic constriction and fitting it with an exponential function $x(t)=a+b\exp(\dot{\varepsilon}t)$.
The cluster size distributions before and after the constriction were evaluated from image sets containing between 1000 and 10000 objects (cells and clusters) for each experiment sequence, using a combination of classical image processing and a custom convolutional neural network (CNN) algorithm described in puthumana2024dissociation and based on the TensorFlow framework abadi2016tensorflow . After training the CNN with a carefully assembled set of images, achieving an impressive accuracy of approximately 95%, the algorithm classifies flowing objects into single cells, doublets, triplets, quadruplets, and larger objects, with their respective relative populations denoted as $S_{i}$, $D_{i}$, $T_{i}$, $Q_{i}$, and $N_{i}$ (where $i=0$ before and $i=1$ after the constriction as defined in Fig. 2a). These populations are normalized by the total number of detected RBCs (aggregated or not), ensuring that $S_{i}+2D_{i}+3T_{i}+4Q_{i}=1$. Aggregates of five cells or more are very rare and are not included in this normalization.
Assuming that aggregates of 5 cells or more are negligible ($N_{i}\ll 1$), mass conservation between the initial size distribution before the constriction $(S_{0},D_{0},T_{0},Q_{0})$ and after the constriction $(S_{1},D_{1},T_{1},Q_{1})$ allows to compute dissociation probabilities $P_{d}$ (for doublets), $P_{t}$ (triplet breaking into a doublet and a single cell), $P_{q1}$ and $P_{q2}$ (quadruplet breaking into respectively two doublets or a single cell and a triplet), with $P_{q}=P_{q1}+P_{q2}$, with the additional assumption that $P_{q1}=2P_{q2}$ (assuming that all cell-cell bonds in a quadruplet are equivalent) as done previously puthumana2024dissociation .
Figure 3: Dissociation probabilities of aggregates of 2, 3 and 4 cells (doublets (a), triplets (b) and quadruplets (c)) for RBCs from layers $L_{2}$ to $L_{4}$ obtained by density fractionation and from whole blood (WB) samples (data from puthumana2024dissociation ). Continuous lines represent fitting curves using equation 1. All series represent data from 3 different donors.
III Results and Discussion
[…]
The results are summarized in Fig. 3 in which the dissociation probabilities $P_{d}$, $P_{t}$, $P_{q}$ of aggregates of cells from density layers $L_{2}$ to $L_{4}$ and from whole blood ($WB$) are plotted as a function of extension rate $\dot{\varepsilon}$, with a 3g/dL Dextran buffer as suspending medium.
As previously reported puthumana2024dissociation , all disaggregation probabilities follow a characteristic sigmoid curve when increasing the extension rate, with a strong jump in the $80-200$ s-1 allowing to define a critical extension rate $\dot{\varepsilon}_{c}$ corresponding to the maximum rate of change in dissociation probabilities. To do so, experimental data in Fig. 3 is fitted with a sigmoid curve defined as:
$P(\dot{\varepsilon})=A+B\tanh\left(C(\dot{\varepsilon}-\dot{\varepsilon}_{c})\right)+D\dot{\varepsilon}~{},$ (1)
where $A$, $B$, $C$, $D$ and $\dot{\varepsilon}_{c}$ are fitting parameters. The determination coefficient $R^{2}$ for all data series in Fig. 3 is better than $0.99$, allowing for a reliable determination of $\dot{\varepsilon}_{c}$.
In puthumana2024dissociation , an analysis of the forces acting on cells within a cluster, performed in a reference frame centered on the cluster’s center of mass, concluded that the dissociation dynamics are primarily governed by the balance between the work of viscous drag forces on the cells and the interaction energy between them. This led to the proposal of a scaling law for the critical extension rate required to dissociate a two-cell aggregate (doublet):
$\dot{\varepsilon}_{c}\simeq\frac{\epsilon_{ad}}{12\eta\delta l}~{},$ (2)
which can be generalized for a (linear) aggregate of $N$ cells:
$\dot{\varepsilon}_{c}(N)\simeq\frac{\epsilon_{ad}}{12(N/2)\eta\delta l}~{},$ (3)
[…]
As suggested by the scaling $\dot{\varepsilon}_{c}(N)\sim 1/N$, larger aggregates tend to dissociate at lower extension rates compared to smaller aggregates as larger objects experience higher drag forces in the extensional field, as seen in Fig. 4(a-c) where data for each layer is represented separately. This is also reflected in Fig. 4(d) where dissociation probabilities are plotted as a function of $\eta\dot{\varepsilon}N/2$ as suggested by Eq. 3: data for triplets ($P_{t}$) and quadruplets ($P_{q}$) rather nicely collapse with the data for doublets ($P_{d}$) from $L_{2}$, $L_{3}$ and $WB$. The collapse is not as good for $L_{4}$ (cells with the highest density - corresponding to older cells) for which $\dot{\varepsilon}$ only weakly depends on $N$ as visible in Fig. 4(c) where the lateral shift is smaller and dissociation probabilities for triplets and quadruplets are almost identical. We shall discuss this below.
[…]
Comparing the data from different density layers in Figs. 3 and 4 shows that RBCs from denser layers tend to form more robust aggregates that require higher extensional stresses to dissociate. When increasing $\dot{\varepsilon}$, cells from $L_{2}$ are the first to dissociate, followed by $L_{3}$, and then by $L_{4}$. The experimental critical extension rate $\dot{\varepsilon}_{c}$ defined as the inflection point in dissociation curves of Fig. 3 - 4 is shown in Fig. 5. On average, aggregates from $L_{4}$ require extensional stresses that are approximately twice as high as the value needed for $L_{2}$ to reach a similar dissociation probability (e.g. for quadruplets) with a significant increase (up to 65%) of $\dot{\varepsilon}_{c}$. Note that the data for $WB$ is roughly between the data for $L_{2}$ and $L_{3}$ in Figs 3 and 4(d), which is consistent with the fact that $L_{2}$ corresponds to the peak of the RBC density distribution as seen in Fig. 1 and previous works mohandas88densdistrib ; Ermolinskiy2020Oct ; nouaman2023effect .
The very strong shift in extension rate observed for $L_{4}$ compared to a more modest one between $L_{2}$ and $L_{3}$ is reminiscent of the variation of deformability reported by Nouaman et al. nouaman2023effect who measured deformability indexes (DI) and DI variations with strain rate that are drastically reduced for $L_{4}$ using the same density fractionation, or with earlier ektacytometry measurements that also show more pronounced changes for the densest fractions of RBCs mohandas1989cell . In comparison, optical tweezer measurements Ermolinskiy2020Oct revealed a significant and gradual increase of about 20% of the aggregation force between $L_{2}$ and $L_{4}$ in autologous plasma and about 10% in a 5g/dL Dextran 70 kDa solution. Although the suspending media are different, the amplitude of the measured effect are comparable with the evolution of $\dot{\varepsilon}_{c}$ for doublets in Fig. 5: the critical extension rate for doublets is about 10% higher in $L_{4}$ compared to $L_{2}$. In parallel, the value of $P_{d}$ is also significantly lower in $L_{4}$ at any value of $\dot{\varepsilon}$ (see Fig. 3), reflecting stronger aggregability. The aggregate sizes of $L_{4}$ cells have been shown to be smaller compared to other layers in cytometry measurements of Percoll separated samples huang2011human . Following the analysis leading to Equation 3, aggregates of older cells may therefore require a higher extension rate to experience the same total drag force as slightly bigger young cell aggregates.
The evolution with cell density is however much more contrasted with aggregates of 3 and 4 cells that are much more robust in $L_{4}$. This is also consistent with the fact that in this density layer, aggregates of different sizes (doublets, triplets, quadruplets) dissociate at about the same extension rate and do not obey the scaling with $N$ proposed in equation.3. This can be related to morphological changes that can be observed and have been documented for denser and older cells:
  • cell aging and density increase is associated to dehydration, membrane loss and other alterations that can lead to smaller, less deformable and possibly more spherical cells. A higher occurence of echinocytes in older cells has also been reported huang2011human . These alterations at the cell level have consequences on aggregate morphology.
  • for the same reason as mentioned above, the higher compacity or smaller volume of aggregates of denser cells means that the scaling with $N$ in Eq. 3 is less applicable, as it is based on the assumption that aggregates are linear rouleaux.
  • differences in membrane properties such as surface charge are also likely to contribute hadengue1998erythrocyte in the variations observed in between younger (lighter) and older (denser) cell aggregates.
IV Conclusions
We investigated the dissociation of red blood cell aggregates under extension, focusing on the effects of cell aging, using cell density as an indicator of cell age. A significant variation in the dissociation of aggregates with cell density was observed, both in terms of critical hydrodynamic stress and dissociation probability. This finding is consistent with previous studies on the differences in aggregation strengths between cells of varying ages and confirms that RBC aggregability increases with cell density and age.
In contrast to other aggregability measurement methods, the microfluidic technique we used here provides hydrodynamic conditions that are relevant to flow in the microcirculation, particularly in the bifurcations of capillary networks. Our findings therefore contribute to the understanding of aggregation dynamics in microcirculation and the consequences of cell aging. While it is well known that cell aging in vivo leads to decreased cell deformability, which is the mechanism by which older cells are filtered and removed from the blood in the spleen Moreau2023 , the increased RBC aggregability and aggregate robustness that we quantified in this work are likely to have significant consequences at the microcirculatory level (increased resistance to flow, increased clogging) that may be more severe than the simple decrease in deformability for single cells. This further underscores the physiological importance of removing aging cells from circulation.
Altered properties through which aggregability and robustness of aggregates may be increased for denser and older cells were mentioned, namely RBC shape, volume and surface changes, decreased deformability and altered membrane surface properties. Although these effects are not dissociable in our study, we suggest that the same technique can be used to study the influence of each parameter separately, for instance after modifying RBC deformability in a controlled way using diamide or glutaraldehyde, or altering RBC volume by suspending them in hyper or hypo-osmotic media.
Acknowledgements.
The authors thank M. Van Melle-Gateau (LIPhy, CNRS-UGA) for technical assistance (microfabrication), T. John, M. Nouaman, A. Darras and S. Reckenwald for experimental advice; M. P M and T.P. thank G. Ghigliotti, E. Franceschini and G. Coupier for discussions and suggestions. This work was supported by CNES (Centre National d’Etudes Spatiales, DAR ID 8106 ”Rhéologie sanguine”), and ANR project ANR-15-IDEX-02 (IDEX Université Grenoble Alpes, ”Aide à la mobilité internationale des doctorants”).